How fast will an 800 kg car accelerate if it is pushed with 4000N of force?
1
Expert's answer
2020-11-25T10:52:17-0500

By the definition of the Newton's Second Law of Motion, we have:


"F=ma,""a=\dfrac{F}{m}=\dfrac{4000\ N}{800\ kg}=5\ \dfrac{m}{s^2}."

Answer:

"a=5\ \dfrac{m}{s^2}."